OpenStudy (anonymous):

Solve the system using elimination. x-2y=5 2x-4y=-1

OpenStudy (zpupster):

x-2y=5 2x-4y=-1 mult the 1st equation by -2 -2(x-2y=5) -2x+4y=-10 then add to 2nd equation 2x-4y=-1 the variables add to zero leaving there is no solution to this system of equations

OpenStudy (anonymous):

Thank you, I thought for sure I was doing something wrong.